OpenStudy (anonymous):

What are the coordinates of the midpoint of the segment whose endpoints are J(3, –1) and K(9, 5)? (1, 2) (3, 2) (3, 3) (6, 2)

OpenStudy (anonymous):

Add each component to the other, then divide by 2.

OpenStudy (igreen):

Use the midpoint formula. \((\dfrac{x_1 + x_2}{2}, \dfrac{y_1+y_2}{2})\)

OpenStudy (anonymous):

J+K=(12, 4)

OpenStudy (anonymous):

D?

OpenStudy (igreen):

Yep, you got it.